OpenStudy (anonymous):

Neeed some help :) thanks 33.4.5z-2.5z=24 22.9c-5=13

OpenStudy (anonymous):

that doesn't work. :/

OpenStudy (anonymous):

33.4.5z here there are two decimals ..its not possible

OpenStudy (anonymous):

9c-5=13 w/7-2=9 4.5z-2.5z=24 are the 3 problems I neeed some help on. THANK YOU

OpenStudy (anonymous):

w/7 -2 = 9 -> w/7 = 2 -> w = 14 this is wrong w/7 =9+2 =11 w= 11* 7= 77 ans

OpenStudy (anonymous):

w/7 -2 = 9 -> w/7 = 2 -> w = 14 is wrong

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

:)